OpenStudy (anonymous):

Which of the following is the solution set of 18 - 3n + 2 = n + 20 - 4n?

OpenStudy (anonymous):

what are the solutions ?

OpenStudy (anonymous):

not sure

OpenStudy (anonymous):

there aren't any answer choices ?

OpenStudy (anonymous):

ya my fault -14m - 7m + 1 = -7m + 1 -14m + 7m + 1 = -7m + 2 14m - 7m - 1 = -7m + 1